OpenStudy (anonymous):

if it takes 12 people 8 hours to do a certain job, how many hours would it take 3 people to do 1/4 of the same job?

OpenStudy (anonymous):

8 hours

OpenStudy (cwrw238):

find how long it takes 3 people to do the whole job then divide that by 4

OpenStudy (anonymous):

@cwrw238 If Julian is correct, you would need to multiply by 4 instead of dividing.

OpenStudy (cwrw238):

for whole job time by 3 people = 12 * 8 ----- = 32 hours 3 - now divide that by 4...

OpenStudy (anonymous):

But why not 12/8=3/x as a proportion?

OpenStudy (cwrw238):

no - the number of hours for 3 people must take more time than when 12 people do it ( time 4)

OpenStudy (anonymous):

I see.

OpenStudy (cwrw238):

this is inverse proportionality less people - more hours

OpenStudy (cwrw238):

h = k/p when h = 8, p = 12 8 = k/12 k = 96 so h = 96/p when p = 3 h = 96/3 = 32